# A citation index bridging Hirsch’s $h$ and Egghe’s $g$

### v1

@@ 3. An alternative citation index $\nu$
The citation indexes $h$, $\nu$, $\bar{\nu}$, $g$ and $g^{*}$ are in the following relations:
$h$ $\leq\nu\leq g^{*},$ (13)
$h$ $\leq\bar{\nu}\leq g.$ (14)
@@
We only prove inequalities (13); the proof of (14) is similar. First, by definition of $h$ in (3) we can write
$h\leq\sum_{i=1}^{m}{\bf 1}_{\{x_{i}\geq h\}}=\frac{1}{h}\sum_{i=1}^{m}x_{i}\mbox{$\;\!$}{\bf 1}_{\{x_{i}\geq h\}}.$
@@
Next, let the maximizing sum in (11) be expressed as
$\nu^{2}\leq\sum_{i=1}^{m}x_{i}\mbox{$\;\!$}{\bf 1}_{\{x_{i}\geq\nu\}}=\sum_{i=1}^{k^{*}}x_{i},$ (15)
thus involving $k^{*}$ terms $x_{i}$, each one satisfying the inequality
$x_{i}\geq\nu\qquad(i=1,\dots,k^{*}).$ (16)
If $k^{*}\leq\nu$ then from (15) we obtain
$\nu^{2}\leq\sum_{i=1}^{k^{*}}x_{i}\leq\sum_{i=1}^{\nu}x_{i},$
and it follows from definition (10) that $g^{*}\geq\nu$. Alternatively, if $k^{*}\geq\nu$ then, using (16), we can write
$\sum_{i=1}^{k^{*}}x_{i}\geq\sum_{i=1}^{\nu}x_{i}\geq\sum_{i=1}^{\nu}\nu=\nu^{2},$
@@
Table 1: Examples
@@
4. Parametric family $(\nu_{\alpha})$
@@
$\nu_{\alpha}\equiv\nu_{\alpha}(\bm{x})=\max\left\{j\geq 1\colon\sum_{i=1}^{m}x_{i}^{\alpha}\mbox{$\;\!$}{\bf 1}_{\{x_{i}\geq j\}}\geq j^{\alpha+1}\right\}\!.$ (17)
@@
$h=\nu_{0},\qquad\nu=\nu_{1}.$ (18)
Like in (3) and (11), the existence and uniqueness of the maximum $\nu_{\alpha}$ is self-evident, noting that the sum is a decreasing function of $j$ while the right-hand side is strictly increasing. It is also straightforward to verify that $\nu_{\alpha}$ satisfies (C1)–(C3). We also observe the monotonicity of the family $(\nu_{\alpha})$.
Lemma 5.
@@
$\nu_{\alpha}=\max\left\{j\geq 1\colon\sum_{i=1}^{m}\left(\frac{x_{i}}{j}\right)^{\mbox{$\>\!\!$}\alpha}{\bf 1}_{\{x_{i}\geq j\}}\geq j\right\}\!,$ (19)
@@
It is also interesting to understand the meaning of the limiting value $\nu_{\infty}$.
Lemma 6.
For a citation vector $\bm{x}=(x_{1},\dots,x_{m})$, denote by $\ell_{1}=\sum_{i=1}^{m}{\bf 1}_{\{x_{i}=x_{1}\}}$ the multiplicity of the top citation number $x_{1}$. Then
$\nu_{\infty}(\bm{x})=\begin{cases}x_{1}-1&\text{if }\,\ell_{1}<x_{1},\\ x_{1}&{\text{otherwise}.}\end{cases}$ (20)
@@
Easily follows using (19) by noting that $(x_{1}/j)^{\infty}$ equals $\infty$, $1$ or 00 according as $j<x_{1}$, $j=x_{1}$ or $j>x_{1}$, respectively. ∎
5. Discussion and conclusions
We have introduced some new citation indexes starting with $\nu=\nu_{1}$, and investigated their relations with the classical indexes $h$ and $g$.
@@ 5. Discussion and conclusions
As already mentioned, the $h$-index is straightforward and informative, but it is limited by only acknowledging the fact of a high citation but not the actual number of citations. In contrast, the $g$-index is based exclusively on the citations of a few top papers, but ignoring the “footing” of lower-cited papers.
@@
Our synthetic proposal of the $\nu$-index is designed so as to take into account both higher and lower cited papers, which may assess the individual’s productivity in a more fair and balanced way. Indeed, we have seen that the $\nu$-index is in a sense bridging Hirsch’s $h$ and Egghe’s $g$.
Furthermore, the spectrum of the indexes $(\nu_{\alpha})$ provides a flexible toolkit that allows one either to enhance or to inhibit the input from top-cited papers, as required.
@@
Of course, it goes without saying that none of these, or any other indexes known in the literature is perfect and should replace the rest. In fact, a reasonable practical recommendation may be to choose a few indexes to judge someone’s academic achievement, depending on the assessment requirements and also on the specific features of the scientific domain.
In this regard, it may be useful to choose the parameter $\alpha$ in the index $\nu_{\alpha}$ according to certain individual features of the citation vector $\bm{x}$, in the spirit of limit theorems for norms of random vectors (Schlather, 2001; Bogachev, 2006). We will address this issue in our future work.
As another direction of research, it is important to develop methods of evaluation of a preferred citation index under a suitable statistical model of citation frequencies. We advocate (Nuermaimaiti et al., 2021; Bogachev et al., 2023; Nuermaimaiti, 2023) the use of the so-called limit shape of the random citation profile, which enables one to use an appropriate rescaling of the “tail” of the frequency distribution to extract visual information about its behavior. The idea of the limit shape stems from theory of random integer partitions (Vershik, 1996) and was first applied to the citation context by Yong (2014).
To give just one example of utility of the limit shape, consider a power law frequency model $f_{j}=j^{-a}\mbox{$\>\!\!$}/\zeta(a)$ ($j\geq 1$), where $\zeta(a)=\sum_{j\geq 1}j^{-a}$ is the Riemann zeta function ($a>1$). Egghe & Rousseau (2006) proposed the following formula for the $h$-index,
$h\approx m^{1/a},$ (21)
where $m$ is the number of published papers. This expression easily follows upon the replacement of sums emerging in the power law model by the corresponding integrals, which can be explicitly computed. As an alternative, we argue that, even though the power law is scale free (in that $A^{a}f_{\mbox{$\;\!\!$}Ax}=f_{x}$, for any scaling factor $A>0$), the limit shape exists and is given by the formula $y=x^{1-a}/c_{a}$, where $c_{a}=(a-1)\mbox{$\;\!$}\zeta(a)>1$. As a result, the $h$-index can be estimated as follows (Nuermaimaiti, 2023),
$h\approx\left(\frac{m}{c_{a}}\right)^{\mbox{$\>\!\!$}1/a}\!.$ (22)
Estimator (22) is more conservative compared to (21), but most importantly, it retains the number-theoretic flavor of the original discrete distribution (through the term $\zeta(a)$).

### v2

@@ 3.2 Checking the basic properties
3.3 Main result – ordering relations between the indexes
@@
The citation indexes $h$, $\nu$, $\bar{\nu}$, $g$ and $g^{*}\!$ are in the following ordering relations:
[…]
@@
We only prove the inequalities for $\nu$; the proof for $\bar{\nu}$ is similar. First, by definition of $h$ in (3) and (4) we can write
$h\leq m_{*\mbox{$\;\!\!$}}(h)=\sum_{i=1}^{m}\bm{1}_{\{x_{i}\geq h\}}\leq\frac{1}{h}\sum_{i=1}^{m}x_{i}\mbox{$\>\!$}\bm{1}_{\{x_{i}\geq h\}}=\frac{1}{h}S_{m_{*\mbox{$\;\!\!$}}(h)}.$
@@
Next, the maximizing sum in (10) is expressed as
$\nu^{2}\mbox{$\>\!\!$}\leq\sum_{i=1}^{m}x_{i}\mbox{$\>\!$}\bm{1}_{\{x_{i}\geq\nu\}}=S_{m_{*\mbox{$\;\!\!$}}(\nu)},$ (14)
thus involving $m_{*\mbox{$\;\!\!$}}(\nu)$ terms $x_{i}$ satisfying the inequality
$x_{i}\geq\nu,\qquad i=1,\dots,m_{*\mbox{$\;\!\!$}}(\nu).$ (15)
If $m_{*\mbox{$\;\!\!$}}(\nu)\leq\nu$ then from (14) we obtain (adding fictitious zeros if $\nu>m$)
$\nu^{2}\mbox{$\>\!\!$}\leq S_{m_{*\mbox{$\;\!\!$}}(\nu)}\leq S_{\nu},$
and it follows from definition (9) that $g^{*}\!\geq\nu$. Alternatively, if $m_{*\mbox{$\;\!\!$}}(\nu)\geq\nu$ then, using (15), we can write
$S_{m_{*\mbox{$\;\!\!$}}(\nu)}\geq S_{\nu}\geq\nu^{2}\mbox{$\>\!\!$}$,
@@
3.4 R code and some simple examples
A simple R code to calculate various indexes is given below:
# indexes h, nu, nu.bar, g, g.star
ind <- function(x) # x = input citation vector
{ x <- sort(x, decreasing = TRUE) # ordering
m <- length(x) # number of papers
# h
h <- 0
while (h < length(x) && x[h + 1] >= h + 1)
{ h <- h + 1
# nu
nu <- 0
while (sum(x[which(x >= (nu + 1))])
>= (nu + 1)^2)
{ nu <- nu + 1
# nu.bar
nu.bar <- min(nu,m)
# g
g <- max(which(cumsum(x) >= (1:m)^2))
# g.star
if (sum(x) >= m^2)
{ g.star <- floor(sqrt(sum(x)))
else
{ g.star <- max ( which (cumsum(x) >= (1:m)^2))
}# Printing the output:
[…]
cat("h =", h, "nu.bar =", nu.bar, "nu =", nu,
"g =", g, "g.star =", g.star)
Example (John Nash case):
x <- c(2000,2000,1500,1000,400,250,100,100)
ind(x)
# x = ( 2000 2000 1500 1000 400 250 100 100 );
# h = 8 nu.bar = 8 nu = 85 g = 8 g.star = 85
@@ 3.4 R code and some simple examples
Table 1: Illustrative examples of different indexes.
@@
3.5 Cases of equality
One observation from Table 1 is that, occasionally, some of the indexes may coincide, which warrants a question of exploring the cases of equalities in (13). The possible equality $h=g^{*}\!$ was addressed by Egghe et al. (2019).
Theorem 2.
The equalities in the index inequalities (13) of Theorem 1 hold if and only if the following conditions are satisfied, respectively:
(a) $h=\nu\colon$ $S_{m_{*\mbox{$\;\!\!$}}(h+1)}<(h+1)^{2};$
(b) $\nu=g^{*}\!{\colon}$ $S_{\nu+1}<(\nu+1)^{2};$
(c) $h=\bar{\nu}\colon$ $h=m$, or $h<m$ and $S_{m_{*\mbox{$\;\!\!$}(h+1)}}<(h+1)^{2};$
(d) $\nu=g\colon$ $\nu=m$, or $\nu<m$ and $S_{\nu+1}<(\nu+1)^{2}.$
Proof.
Since it is always true that $\nu\geq h$ (see (13)), the equality $\nu=h$ simply means that $\nu<h+1$. But, according to definition (11), the latter inequality is equivalent to $S_{m_{*\mbox{$\;\!\!$}}(h+1)}<(h+1)^{2}$, which is the claim of part (a). Essentially the same argument proves part (c), except that, due to the bound $\bar{\nu}\leq m$, a special case arises if $h=m$, which automatically implies $\bar{\nu}=m$.
Similarly, due to (13) we have $g^{*}\!\geq\nu$, while the inequality $g^{*}\!<\nu+1$ is equivalent to $S_{\nu+1}<(\nu+1)^{2}$, according to definition (9), and the claim of part (b) follows. The same argument applies to part (d), with an additional consideration of the special case $\nu=m$. ∎
[…]
3.6 Data example
Here, we illustrate the calculation of the various citation indexes for real data collected by the first-named author (available online at https://github.com/Ruheyan/WoS-citation-data/tree/main). [^1] Similar citation data was used in a conference paper (Nuermaimaiti et al., 2021) and PhD thesis (Nuermaimaiti, 2023). The dataset comprises citation counts, with a cut-off date of 19th September 2022, of 3,615 papers (with 73,730 citations in total) of 111 authors who published a paper in the first 10 issues of Electronic Journal of Probability (EJP), vol. 24 (2019) (https://projecteuclid.org/journals/electronic-journal-of-probability/volume-24/issue-none). The data were derived from the Web of Science (WoS, 2025).
Figure 1: Index plots for the EJP dataset, showing triplets of indexes ($h\leq\nu\leq g^{*}$ or $h\leq\bar{\nu}\leq g$) normalized by the number of published papers per author. The authors are ranked in increasing order with respect to the parameter $h/m$.
Fig. 1 shows the plots representing the indexes $h$, $\bar{\nu}$, $\nu$, $g$, and $g^{*}$ (in triplets, for ease of comparison) for all 111 authors, normalized by the number of papers per author. The calculated values confirm the inequalities of Theorem 1, but one can observe that the new index ($\nu$ or $\bar{\nu}$) tends to be closer to the upper bound $g^{*}\!$ or $g$, respectively. Furthermore, Table 2 presents correlations between different indexes — not surprisingly, they are all strongly positively correlated (especially in the “sister” pairs $(\nu,\bar{\nu})$ and $(g,g^{*})$), but correlation with the number of papers ($m$) is weaker.
Table 2: Pairwise correlations across the citation indexes and the number of published papers ($m$).
index $h$ $\nu$ $\bar{\nu}$ $g$ $g^{*}$ $m$
$h$ 1.0000 0.9649 0.9646 0.9656 0.9725 0.8044
$\nu$ 0.9649 1.0000 0.9998 0.9932 0.9978 0.7743
$\bar{\nu}$ 0.9646 0.9998 1.0000 0.9942 0.9978 0.7768
$g$ 0.9656 0.9932 0.9942 1.0000 0.9967 0.8046
$g^{*}$ 0.9725 0.9978 0.9978 0.9967 1.0000 0.7893
$m$ 0.8044 0.7743 0.7768 0.8046 0.7893 1.0000
4 Parametric family $(\nu_{\alpha})$
4.1 Definition and monotonicity
@@
$\nu_{\alpha}\equiv\nu_{\alpha}(\bm{x})=\max\left\{j\geq 1\colon\sum_{i=1}^{m}x_{i}^{\alpha}\mbox{$\>\!$}\bm{1}_{\{x_{i}\geq j\}}\geq j^{\alpha+1}\right\}\!.$ (16)
@@
$\nu_{0}=h,\qquad\nu_{1}=\nu.$
Like in (3) and (10), the existence and uniqueness of the maximum in (16) is self-evident, noting that the sum is a decreasing function of $j$ while the right-hand side is strictly increasing. It is straightforward to verify that $\nu_{\alpha}$ satisfies (C1)–(C3). We also observe the monotonicity of the family $(\nu_{\alpha})$.
Theorem 3.
@@
$\nu_{\alpha}=\max\left\{j\geq 1\colon\sum_{i=1}^{m}\left(\frac{x_{i}}{j}\right)^{\mbox{$\>\!\!$}\alpha}\!\bm{1}_{\{x_{i}\geq j\}}\geq j\right\}\!,$ (17)
@@
As an illustration of sensitivity and fluidity of $\nu_{\alpha}$, in the John Nash case it is easy to check that, for example, for $\alpha=0.5$ we have $\nu_{0.5}=35$, compared to $\nu_{1}\mbox{$\>\!\!$}=g^{*}\!=85$. R code to calculate $\nu_{\alpha}$ is given below:
# nu.alpha
x <- sort(x, decreasing=TRUE)
nu.alpha <- function(alpha)
{ sapply(alpha, function(a)
{ nu <- 0
while (sum((x[x >= (nu + 1)] / (nu + 1))^a)
>= (nu + 1))
{ nu <- nu + 1
return(nu)
# Plotting the output:
curve(nu.alpha, col = "red", lwd = 2,
xlim = c(0, max(x)+20), ylim = c(1, max(x)),
xlab = expression(paste(alpha)),
ylab = expression(paste(nu[alpha])),
main = bquote(paste(bold("x "), "= (",
.(toString(x)), ")")))
The next Fig. 2 illustrates the behavior of the function $\nu_{\alpha}$ for some examples from Table 1. The reader may also find it interesting to run this code on the citation data of John Nash.
Figure 2: Illustrative graphs of the index $\nu_{\alpha}$ as a function of parameter $\alpha\in[0,\infty)$. Note the values $\nu_{0}=h$, $\nu_{1}=\nu$, and $\nu_{\infty}=x_{1}-1$ (cf. (18)).
4.2 The limit as $\alpha\to\infty$
It is interesting to understand the meaning of the limiting value $\nu_{\infty}=\lim_{\alpha\to\infty}\nu_{\alpha}$.
Theorem 4.
For a citation vector $\bm{x}=(x_{1},\dots,x_{m})$, denote by $\ell_{1}=\sum_{i=1}^{m}\bm{1}_{\{x_{i}=x_{1}\}}\equiv m_{*\mbox{$\;\!\!$}}(x_{1})$ the multiplicity of the top citation $x_{1}=\max\{x_{i},\,1\leq i\leq m\}$. Then
$\nu_{\infty}(\bm{x})=\begin{cases}x_{1}-1&\text{if }\,\ell_{1}<x_{1},\\ x_{1}&\text{if }\,\ell_{1}\geq x_{1}.\end{cases}$ (18)
@@ 4.2 The limit as $\alpha\to\infty$
Follows using (17) by noting that $(x_{1}/j)^{\infty}$ equals $\infty$, $1$ or 0 according as $j<x_{1}$, $j=x_{1}$ or $j>x_{1}$, respectively. ∎
5 Conclusion
@@
We have introduced some new citation indexes starting with $\nu=\nu_{1}$, and investigated their relations with the classical indexes $h$ and $g$. As already mentioned, the $h$-index is straightforward and informative, but it is limited by only acknowledging the fact of a high citation but not the actual number of citations. In contrast, the $g$-index is based exclusively on the citations of a few top papers, but ignoring the “footing” of lower-cited papers.
@@
Our synthetic proposal of the $\nu$-index is designed so as to take into account both higher and lower cited papers, which may assess the individual’s productivity in a more fair and balanced way. Indeed, we have seen that the $\nu$-index is in a sense bridging Hirsch’s $h$ and Egghe’s $g$. Furthermore, the spectrum of the indexes $(\nu_{\alpha})$ provides a flexible toolkit that allows one either to enhance or to inhibit the input from top-cited papers, as required.
@@
Of course, it goes without saying that none of these, or any other indexes known in the literature, is perfect and should replace the rest. In fact, a reasonable practical recommendation may be to choose a few indexes to judge someone’s academic achievement, depending on the assessment requirements and also on the specific features of the scientific domain. In this regard, it may be useful to choose the parameter $\alpha$ in the index $\nu_{\alpha}$ according to certain individual features of the citation vector $\bm{x}$, in the spirit of limit theorems for norms of random vectors (Bogachev, 2006; Schlather, 2001). We will address this issue in our future work.
In conclusion, we reiterate that prudence, maturity and care should be exercised when using citation indexes in social practice, especially making sure to avoid misuse and/or abuse of their utility as predictors of future performance and productivity. Although citation indexes succinctly grasp some objective aggregated information from citation records, they are deceptively easy to compute, replacing individual research track records with a simple number, while these results should be verified and complemented by human evaluation by experts.
The scientometrics community has quickly realized, and extensively documented, the growing threat of misusing the $h$-index and other indicators for far reaching and often unjustified implications in the social interpretation (see, e.g., Alonso et al., 2009; Costas & Bordons, 2007; Hicks et al., 2015; Thelwall, 2025; Waltman, 2016, and further references therein). These concerns and wide discussions have led to the creation and promotion of good practice protocols, such as the San Francisco Declaration on Research Assessment (DORA, 2020) or the Leiden Manifesto (Hicks et al., 2015).
The risks are further amplified by the fast growing use of Artificial Intelligence (AI) including Large Language Models such as ChatGPT, whereby the responsibility for conclusions and extrapolations may be delegated inadvertently to the computer (Thelwall, 2025). Although deployment of AI for assistance in technical analyses and summarization is an inevitable and welcome trend, the best vaccine against misuse and abuse is to combine formal calculations and summaries with a robust comparison against the specific domain “golden standards”, based on an objective expert evaluation and enhanced by a reproducible and unbiased statistical analysis.
Funding
R.N. acknowledges funding from the School of Mathematics at the University of Leeds for attending the STI-ENID 2025 conference to present a poster based on this work.
Author contributions
L.B. and R.N. designed the study and wrote the first draft of the paper. L.B. and J.V. contributed to the mathematical derivation and interpretation of the results. All authors agreed with the final version of the paper.
Previous presentation
These results were presented as a poster at the 29th Annual International Conference on Science and Technology Indicators (STI-ENID 2025), Bristol, UK, 3–5 September 2025 (https://virtual.oxfordabstracts.com/event/74626/submission/143).
Data availability
EJP dataset used in this article is available online at https://github.com/Ruheyan/WoS- citation-data/tree/main.